OpenStudy (rational):

\(f(x)\) is a polynomial with integer coefficients. show that \[\frac{f(b)-f(a)}{b-a}\] is always an integer for all distinct integers \(a,b\)

OpenStudy (freckles):

say \[f(x)=m_0+m_1x+m_2x^2+ \cdots + m_nx^n \\ \text{ so we want \to show there exist integer } k \text{ such that } \\ (b-a)|[(m_0+m_1b+m_2b^2+\cdots m_nb^n)-(m_0+m_1a+m_2a^2+ \cdots +m_nb^n)] \\ \text{ so we have we want \to show } \\ (b-a)|[m_1(b-a)+m_2(b^2-a^2)+m(b^3-a^3)+\cdots m_n(b^n-a^n)]\] but all of those will have a factor (b-a) in common

OpenStudy (kainui):

\[f(a) \equiv f(b) \mod b-a\] substitute \(k=b-a\) \[f(a) \equiv f(k+a) \mod k\] Since f(k+a) is just a bunch of integers multiplying binomials, all the higher terms containing k in the expansion cancel out and all you're left with is the f(a) terms. Done!

OpenStudy (freckles):

\[ (b-a) \\ (b^2-a^2)=(b-a)(b+a) \\ (b^3-a^3)=(b-a)(b^2+ab+a^2) \\ \cdots \\ (b^n-a^n)=(b-a)(b^{n-1}+ab^{n-2}+ \cdots +a^{n-1}) \]

OpenStudy (freckles):

\[m_1(b-a)+m_2(b^2-a^2)+\cdots m_n(b^n-a^n) \\ =(b-a)(m_1+m_2(b+a)+\cdots +m_n(b^{n-1}+ab^{n-2}+ \cdots +a^{n-1}))\]

OpenStudy (freckles):

where that big ugly thing next to the (b-a) is an integer which i called k above

OpenStudy (kainui):

Ahhh yeah that's a good too! There's something kind of fascinating about this, that you can always factor out a \(b-a\) out of \(b^n-a^n\) but \(b^n+a^n\) will never be \(c^n\) hahaha.

OpenStudy (kainui):

(for n>2 of course haha)

OpenStudy (rational):

Both are very neat xD Was stuck at this for a while while going through the solution for a related problem http://math.stackexchange.com/questions/83592/prove-that-fx-has-no-integer-roots

OpenStudy (kainui):

Oh I don't know what does this mean \[\mathbb{Z}[x]\]